@ See page 78 05 Question (2 points) In addition to finding the optimal bundles given prices and income, utility maximizationd
Answer #1

u(x1 , x2) = min{48x1 , 24x2}  

budget constraint

p1x1 + p2x2 = m (i)

optimal occurs at kink 48x1 = 24x2  

2x1 = x2

Now substitute   x2 =   2x1 in equation   (i)

p1x1 + p2x2 = m

p1x1 + p2( 2x1 ) = m

x1(p1 + 2p2) = m

x1 = m/(p1 + 2p2)

x*1(p1,p2,m) =    m/(p1 + 2p2)  

x2 = 2m/(p1 + 2p2)  

x*2(p1,p2,m) =    2m/(p1 + 2p2)